If the equation (1-q+(p^2)/2)x^2+p(1+q)x+q(q-1) +((p^2)/2)=0 has two equal roots, prove that p^2=4q


1
Expert's answer
2022-02-28T09:05:05-0500

Quadratic equation has two equal roots if


"1-q+\dfrac{p^2}{2}\not=0"

"D=(p(1+q))^2-4(1-q+\dfrac{p^2}{2})(q(q-1)+\dfrac{p^2}{2})=0"

"p^2(1+q)^2-(2-2q+p^2)(2q^2-2q+p^2)=0"

"p^2(1+2q+q^2)-(2-2q)(2q^2-2q)-p^2(2-2q)"

"-p^2(2q^2-2q+p^2)=0"

"-p^4+p^2(1+2q+q^2-2+2q-2q^2+2q)"

"-(2-2q)(2q^2-2q)=0"

"p^4+p^2(q^2-6q+1)-4q(q-1)^2=0"

If "p^2=4q"


"16q^2+4q^3-24q^2+4q-4q^3+8q^2-4q=0,"

True for "q\in \R."

Or


"p^2=\dfrac{-4q(q-1)^2}{4q}=-(q-1)^2"

The only solution is "p^2=0" for "q=1."

But in this case


"1-q+\dfrac{p^2}{2}=1-1+0=0"

and we have an equation


"0=0"

Hence the quadratic equation


"p^4+p^2(q^2-6q+1)-4q(q-1)^2=0"

has the only solution which satisfies our task


"p^2=4q."

In this case the quadratic equation


"(1-q+\dfrac{p^2}{2})x^2+p(1+q)x+q(q-1)+\dfrac{p^2}{2}=0"

has two equal roots.